OpenStudy (perii224):

QUICK QUESTION WILL MEDAL If a movie is 146 minutes, what time would it end if it started at 11:20 and what time would it end if it started at 12:20?

OpenStudy (thatonegirl_):

well an easy way to solve these is to figure out how many hours and then how many minutes are left. So 146-60=86 (1 hour 86 minutes) 86-60=26 (total is 2 hours and 26 minutes). So add this to the start time, 11:20, and you get 1:46

OpenStudy (perii224):

Thank you guys so much! :)